OpenStudy (anonymous):

If y varies directly as x and y is 400 when x is r and y is r when x is 4, what is the numeric constant of variation in this relation?

OpenStudy (anonymous):

First, you need the change the given equation into standard form. 2x+7y=-15 7y = -2x - 15 y = (-2/7)x - 15/7 The linear equation you are finding has a form like this also y = ax + b In order to be perpendicular, the a of this equation you are finding and the a in the given equation must multiply together and give you - 1, so you have (-2/7)a = -1 a = 7/2 So now you have the a, let's plug it in and have y = (7/2)x + b Now you have the point (-1/2,-2) on this line, -1/2 is x and -2 is y, so you plug it in -2 = (7/2)*(-1/2) + b -2 = (-7/4) + b -2 + 7/4 = b -1/4 = b TA-DAH Your answer is y = (7/2)x - 1/4

OpenStudy (anonymous):

y=kx when x=r,y=400 400=kr ...(1) when x=4,y=r r=4k from (1) 400=k*4k k^2=100 k=10,-10 whenk=10,r=4*10=40 when k=-10,r=4*-10=-40

OpenStudy (anonymous):

thank you @surjithayer i am on an exam and couldnt think of the answer. lol, last question, and couldnt think of it